# Evolution of Concepts in Language Model Pre-Training

### v1

@@ A.1 Selection of Activation Function and Regularization Function
The sparsity of natural features is the fundamental hypothesis underlying superposition (Olshausen & Field, 1997; Elhage et al., 2022; Huben et al., 2024). To obtain crosscoder features with optimal sparsity for ideal interpretability and monosemanticity while maintaining reconstruction fidelity, we carefully select the activation function $\sigma(\cdot)$ and the regularization function $\Omega(\cdot)$.
Previous SAE studies predominantly use ReLU activation with L1 regularization (Bricken et al., 2023; Huben et al., 2024). However, this configuration produces weak feature activations that are largely noise, compromising both interpretability and sparsity. We address this by adopting JumpReLU (Rajamanoharan et al., 2024) as the activation function, which eliminates activations below learned thresholds (trained via straight-through estimation (Bengio et al., 2013)).
@@
For regularization, we employ a combination of tanh and quadratic frequency penalties (Smith et al., 2025): the tanh component provides a superior L0 approximation by reducing penalties on strong activations, while the quadratic frequency penalty suppresses high-frequency features. This yields the following batched sparsity loss:
@@ A.3 Distributed Training Strategy for Crosscoders
To efficiently train crosscoders, we employ a head parallelism distributed training strategy, a variant of tensor parallelism (Shoeybi et al., 2019). With $k$ processes where $k$ divides $n_{\text{snapshots}}$, each process handles encoding and decoding for $\frac{n_{\text{snapshots}}}{k}$ source snapshots. Pre-activations are computed via All-Reduce operations (Sanders et al., 2019). Unlike standard tensor parallelism, our approach processes activations from each snapshot separately on individual processes, reducing I/O overhead for reading activations from disk.
@@ A.4 Experiments
All crosscoders are trained on 800M tokens from the SlimPajama corpus using the Adam optimizer (Kingma & Ba, 2017) with $\beta$ values of $(0.9,0.999)$. To prevent straight-through estimation from causing rapid threshold increases, we apply a reduced learning rate to JumpReLU threshold updates via a multiplier on the global learning rate. JumpReLU thresholds are initialized to $0.1$ for all features. The learning rate schedule includes 10% warm-up steps followed by 20% decay steps. We initialize encoders as transposes of their corresponding decoders, with identical initialization matrices across all snapshots. We employ initialization search to identify optimal decoder norms that minimize loss at initialization. Additional hyperparameters are listed in Table 1.
@@
A.5 Comparison to SAEs
To examine whether crosscoders can extract cross-snapshot features and align identical features in consistent directions within the feature space, we compare crosscoder performance against corresponding SAEs trained individually on each Pythia snapshot.
We train SAEs using identical settings and hyperparameters as crosscoders on each snapshot. Figures 8(a) and 8(b) demonstrate that crosscoders achieve comparable performance in L0 norm and explained variance at each snapshot, even when SAEs are optimized for individual snapshots where they should have a theoretical advantage.
A.6 Potential Failure Modes of Crosscoder Feature Alignment
There are two possible failure modes of cross-snapshot crosscoders:
Will crosscoders misalign unrelated features?
Such misalignment would strongly contradict the optimization objective of crosscoders. The distinct activation patterns of unrelated features would conflict when forced into the same dimension. Once a feature activates, it would cause others to activate simultaneously, introducing noise in the reconstruction. In terms of results, misalignment would lead to the polysemanticity of crosscoder features, damaging the consistency of their interpretation, which our interpretability evaluation demonstrates does not occur.
Will crosscoders split shared features?
Suppose features from different snapshots represent the same underlying concept with identical activation patterns, splitting them across multiple dimensions would be suboptimal: it wastes representational capacity in feature space, as both previous works (Gao et al., 2025; Templeton et al., 2024) and Appendix A.4 prove that feature space dimensionality is crucial for better reconstruction fidelity. Nevertheless, in rare cases, we do observe feature splitting among highly active features. These features share semantic and directional similarity but exhibit subtly different activation patterns that justify separate dimensions. We detail this phenomenon in Appendix F.
A.7 Results on Other Layers
@@
To demonstrate feature evolution across all layers, we apply the same visualization strategy from Section 4, plotting decoder norms for 50 randomly selected features per layer (Figures 11 and 12). Most middle layers exhibit the same evolutionary patterns as Layer 6, supporting our findings in Section 4. We would like to note that the majority of features in Layers 0 and 1 exist from initialization, which aligns with the observation that early layers implements more low-level, especially single-token features (He et al., 2024).
@@ E.1 Formalization of IG Attribution Score
To more accurately estimate the causal effect of features, we employ the integrated gradient (IG) version of the attribution score (Sundararajan et al., 2017; Hanna et al., 2024; Marks et al., 2025). IG attribution computes gradients along an interpolation path between baseline and target activations, providing more robust estimates than single-point gradients.
@@
where $\alpha\in\{0,\frac{1}{N},\ldots,\frac{N-1}{N}\}$ linearly interpolates between the baseline and target feature activations. Consistent with Marks et al. (2025), we use $N=10$ interpolation steps for the IG attribution score.
@@ E.2 Induction Task
Transformers exhibit in-context learning capabilities (Brown et al., 2020; Zhao et al., 2021; Gao et al., 2021) through induction heads (Olsson et al., 2022)—circuits that look back over the sequence for previous instances of the current token (A), identify the subsequent token (B), and predict the same completion will occur again (forming sequences [A][B] … [A] → [B]).
@@ E.3 Subject-Verb Agreement Tasks
Subject-verb agreement tasks evaluate whether models can predict verbs in the appropriate grammatical form based on their subjects. We test four variants: (1) Simple—the verb directly follows the subject, e.g., “The parents are”; (2) Across-RC—a relative clause intervenes between subject and verb, e.g., “The athlete that the managers like does”; (3) Within-RC—both subject and verb appear within the relative clause, e.g., “The athlete that the managers like”; (4) Across-PP—a prepositional phrase separates the subject and verb, e.g., “ The secretaries near the cars have”. We use data provided by Marks et al. (2025), sampling 1000 examples from each variant. We use the verb in the wrong form as the counterpart in attribution patching. Results are shown in Figure 6, 15, 16, and 17.
@@ E.4 Indirect Object Identification Task
Indirect Object Identification (IOI) evaluates whether models can correctly predict the indirect object (IO) based on repeated occurrence of subjects (Wang et al., 2023). In IOI tasks, sentences such as “When Mary and John went to the store, John gave a drink to” should be completed with “Mary.” We generate IOI samples following the same strategy as Wang et al. (2023), using template sentences with random person names. We use the subject name as the corrupted counterpart in attribution patching. Results are shown in Figure 18.
@@ Appendix F Crosscoder Feature Splitting
For example, we identify feature 66688, 53542, and 42307 that all activate on long sequences containing repeated text patterns, contributing to the induction task (Figures 19, 20, and 21). Each feature exists only during non-overlapping continuous periods across snapshots, with a drastic emergence and disappearance.
@@
Appendix G More Examples of Feature

### v2

@@ 8 Conclusion
Reproducibility Statement
All our crosscoders are based on open-source models (Pythia suite) and datasets (SlimPajama) with public accessibility. We provide source code for (1) generating model activations on each snapshot, (2) training crosscoders, and (3) analyzing trained crosscoders at https://github.com/OpenMOSS/Language-Model-SAEs. Detailed instructions for replicating our results are provided in the examples/reproduce_evolution_of_concepts/README.md file. We note that training large crosscoders requires substantial computational resources and disk space, as also listed in the README.md file.
@@ A.1 Selection of Activation Function and Regularization Function
The sparsity of natural features is the fundamental hypothesis underlying superposition (olshausen1997sparsecoding; elhage2022superposition; hubun2024sae). To obtain crosscoder features with optimal sparsity for ideal interpretability and monosemanticity while maintaining reconstruction fidelity, we carefully select the activation function $\sigma(\cdot)$ and the regularization function $\Omega(\cdot)$.
Previous SAE studies predominantly use ReLU activation with L1 regularization (bricken2023monosemanticity; hubun2024sae). However, this configuration produces weak feature activations that are largely noise, compromising both interpretability and sparsity. We address this by adopting JumpReLU (rajamanoharan2024jumprelu) as the activation function, which eliminates activations below learned thresholds (trained via straight-through estimation (bengio2013ste)).
@@
For regularization, we employ a combination of tanh and quadratic frequency penalties (smith2025negativesae): the tanh component provides a superior L0 approximation by reducing penalties on strong activations, while the quadratic frequency penalty suppresses high-frequency features. This yields the following batched sparsity loss:
@@ A.3 Distributed Training Strategy for Crosscoders
To efficiently train crosscoders, we employ a head parallelism distributed training strategy, a variant of tensor parallelism (shoeybi2019megatron). With $k$ processes where $k$ divides $n_{\text{snapshots}}$, each process handles encoding and decoding for $\frac{n_{\text{snapshots}}}{k}$ source snapshots. Pre-activations are computed via All-Reduce operations (sanders2019parallel). Unlike standard tensor parallelism, our approach processes activations from each snapshot separately on individual processes, reducing I/O overhead for reading activations from disk.
@@ A.4 Experiments
All crosscoders are trained on 800M tokens from the SlimPajama corpus using the Adam optimizer (kingma2017adam) with $\beta$ values of $(0.9,0.999)$. To prevent straight-through estimation from causing rapid threshold increases, we apply a reduced learning rate to JumpReLU threshold updates via a multiplier on the global learning rate. JumpReLU thresholds are initialized to $0.1$ for all features. The learning rate schedule includes 10% warm-up steps followed by 20% decay steps. We initialize encoders as transposes of their corresponding decoders, with identical initialization matrices across all snapshots. We employ initialization search to identify optimal decoder norms that minimize loss at initialization. Additional hyperparameters are listed in Table 1.
@@
A.5 Results on Other Layers
@@
To demonstrate feature evolution across all layers, we apply the same visualization strategy from Section 4, plotting decoder norms for 50 randomly selected features per layer (Figures 12 and 13). Most middle layers exhibit the same evolutionary patterns as Layer 6, supporting our findings in Section 4. We would like to note that the majority of features in Layers 0 and 1 exist from initialization, which aligns with the observation that early layers implements more low-level, especially single-token features (he2024llamascope).
@@ E.1 Formalization of IG Attribution Score
To more accurately estimate the causal effect of features, we employ the integrated gradient (IG) version of the attribution score (sundararajan2017attribution; hanna2024faithfulness; marks2025sparsefeaturecircuits). IG attribution computes gradients along an interpolation path between baseline and target activations, providing more robust estimates than single-point gradients.
@@
where $\alpha\in\{0,\frac{1}{N},\ldots,\frac{N-1}{N}\}$ linearly interpolates between the baseline and target feature activations. Consistent with marks2025sparsefeaturecircuits, we use $N=10$ interpolation steps for the IG attribution score.
@@ E.2 Induction Task
Transformers exhibit in-context learning capabilities (brown2020fewshot; zhao2021fewshot; gao2021fewshot) through induction heads (olsson2022induction)—circuits that look back over the sequence for previous instances of the current token (A), identify the subsequent token (B), and predict the same completion will occur again (forming sequences [A][B] … [A] → [B]).
@@ E.3 Subject-Verb Agreement Tasks
Subject-verb agreement tasks evaluate whether models can predict verbs in the appropriate grammatical form based on their subjects. We test four variants: (1) Simple—the verb directly follows the subject, e.g., “The parents are”; (2) Across-RC—a relative clause intervenes between subject and verb, e.g., “The athlete that the managers like does”; (3) Within-RC—both subject and verb appear within the relative clause, e.g., “The athlete that the managers like”; (4) Across-PP—a prepositional phrase separates the subject and verb, e.g., “ The secretaries near the cars have”. We use data provided by marks2025sparsefeaturecircuits, sampling 1000 examples from each variant. We use the verb in the wrong form as the counterpart in attribution patching. Results are shown in Figure 7, 16, 17, and 18.
@@ E.4 Indirect Object Identification Task
Indirect Object Identification (IOI) evaluates whether models can correctly predict the indirect object (IO) based on repeated occurrence of subjects (wang2023ioi). In IOI tasks, sentences such as “When Mary and John went to the store, John gave a drink to” should be completed with “Mary.” We generate IOI samples following the same strategy as wang2023ioi, using template sentences with random person names. We use the subject name as the corrupted counterpart in attribution patching. Results are shown in Figure 19.
@@
E.5 Random Baseline of Ablation
Figure 20: Crosscoder feature attribution on the Across-PP variant of the SVA task, with random baselines.
In Figure 20, we add baselines to Figure 7 by ablating/preserving $k$ random features. The results show that our feature selected by attribution patching is far more effective.
@@ Appendix F Crosscoder Feature Splitting
For example, we identify feature 66688, 53542, and 42307 that all activate on long sequences containing repeated text patterns, contributing to the induction task (Figures 21, 22, and 23). Each feature exists only during non-overlapping continuous periods across snapshots, with a drastic emergence and disappearance.
@@
Appendix G Generalizability Studies
In this section, we include more experiments on training and analyzing crosscoders on different base models/initialization/datasets and different numbers of snapshot selection, to examine the generalizablity of our methods.
G.1 Generalization across Model Families/Seeds
Figure 26: Overview of cross-snapshot feature decoder norm evolution in Pythia-160M Seed1.
Figure 27: Overview of cross-snapshot feature decoder norm evolution in Pythia-160M Seed2.
Figure 28: Overview of cross-snapshot feature decoder norm evolution in Alias run of Stanford CRFM GPT-2.
Figure 29: Crosscoder feature attribution on the Across-PP variant of the SVA task, using a crosscoder with 24,576 features trained on Pythia-160M Seed1.
Figure 30: Crosscoder feature attribution on the Across-PP variant of the SVA task, using a crosscoder with 24,576 features trained on Pythia-160M Seed2.
Figure 31: Crosscoder feature attribution on the Across-PP variant of the SVA task, using a crosscoder with 24,576 features trained on the Alias run of Stanford CRFM GPT-2.
We conduct experiments on two additional Pythia 160M models with different random seeds (Pythia 160M Seed1 and Seed2) (vanderwal2025polypythias) and the Alias run of Stanford CRFM’s GPT-2 model  [^3] Obtained from https://huggingface.co/stanford-crfm/alias-gpt2-small-x21.. All of these models have an activation space of 768 dimensions. We train crosscoders of 24,576 features (32x) on Layer 6 of them, and inspect their feature evolution.
Different initializations (Pythia 160M variants).
We observe highly consistent results across differently initialized versions of Pythia 160M. Decoder norms of initialization features follow identical trajectories, and emergent features begin rising around step 1000, exhibiting a clear two-phase pattern with the same transition point (Figure 26 and 27). Attribution experiments confirm that we successfully identify sparse crosscoder features with significant contributions to downstream tasks (Figure 29 and 30).
Different base model (GPT-2).
For Stanford CRFM’s GPT-2 model, we observe a clear two-phase pattern, with emergent features rising around step 1,000. Upon manual inspection of these features, we still find that initialization features exhibit token-level information, while more complex patterns emerge in emergent features.
However, we find a notable difference: due to absolute positional embeddings, initialization feature norms do not peak at the beginning but rather at later steps (with above-threshold norm at beginning). This difference arises because the absolute positional embeddings initially exhibit large norms (larger than the word embeddings) and dominate the activation norm. Since we normalize all activation sources to the same average activation norm $\sqrt{d_{\text{model}}}$, this large positional contribution substantially reduces the observed norms of initialization features. (Figure 28).
To sum up, the statistical-to-feature-learning phase transition still exists across models, despite small differences. Attribution experiments (Figure 31) again confirm the effectiveness of sparse crosscoder features, with generally similar feature evolution patterns for specific downstream tasks.
Other models with accessible checkpoints.
We also investigated other model families, including BLOOM (bigscience2024bloom) and OLMo (olmo20242olmo2furious). Unfortunately, these models either lack sufficient intermediate checkpoints for fine-grained analysis of pretraining dynamics (BLOOM), or lack sufficiently early checkpoints to capture the two-stage transition (BLOOM and OLMo).
G.2 Generalizability across Datasets
Figure 32: Overview of cross-snapshot feature decoder norm evolution in Pythia-160M, trained on FineWeb-Edu dataset.
We conduct experiments on the original Pythia 160M snapshots, but train and analyze 24,576-feature crosscoders on the FineWeb-Edu dataset (penedo2024fineweb), which consists of educational webpages. Our results (Figure 32) show nearly identical patterns to those observed with Pythia 160M crosscoders trained on SlimPajama, exhibiting very similar feature evolution dynamics.
G.3 Sensitivity to Snapshot Selection
Figure 33: Overview of cross-snapshot feature decoder norm evolution in Pythia-160M, trained on 16 snapshots.
All of our above experiments train crosscoders on 32 pre-training snapshots of the language model. To further examine whether snapshot selection affects the trained crosscoder and the captured feature evolution, we train a 24,576-feature crosscoder on 16 pre-training snapshots of Pythia 160M, downsampled from the original set by selecting every other snapshot. The result (Figure 33) shows that the overall feature evolution exhibits nearly identical trends, with only a reduction in temporal resolution.
Appendix H More Examples of Feature
@@
Appendix I The Use of Large Language Models
The use of Large Language Models as an assistive tool in this paper is limited to the following two aspects:
  1. 1.
    The automated generation of feature complexity scores, as detailed in Section 4.2 and Appendix B.
  2. 2.
    Grammar correction and stylistic refinement of the manuscript.